OpenStudy (anonymous):

PROVE THE IDENTITY cos(3t)= cost(1-4sin^2t)

OpenStudy (anonymous):

@Mertsj

OpenStudy (anonymous):

do you start by changing left to cos(2t+t)?

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

so the left side youd follow the cosine identity? and right side youd keep same?

OpenStudy (anonymous):

@Mertsj what do you do then

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

what?

OpenStudy (mertsj):

Then replace cos(2t) with 1-2sin^2t

OpenStudy (anonymous):

and then?

OpenStudy (anonymous):

sin2t with 2sintcost?

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

then?

OpenStudy (anonymous):

can someone help me with my problem

OpenStudy (mertsj):

@shubham1 Stop with the hijacking of another person's post.

OpenStudy (anonymous):

@shubham this isnt where youu ask for help. you put your question in the question box.

OpenStudy (anonymous):

yeah, dont do that. this is my post.

OpenStudy (anonymous):

@Mertsj sorry, so you were saying?

OpenStudy (anonymous):

how do you simplify the left then

OpenStudy (mertsj):

Simplify and then factor out cos t

OpenStudy (anonymous):

so..

OpenStudy (mary.rojas):

@shamil98 ;D

OpenStudy (mertsj):

Then you are done.

OpenStudy (mary.rojas):

She needs more guidance

OpenStudy (anonymous):

@Mertsj Can you finish my question. Will give medal

OpenStudy (anonymous):

please stop that.

OpenStudy (anonymous):

if i factor out cos t can you show me cuz im a little stuck

OpenStudy (mertsj):

\[\cos (2t+t)=\cos 2t \cos t-\sin 2t \sin \sin t=(1-2\sin ^2t) \cos t-2\sin t \cos t \sin t=\] \[\cos t-2\sin ^2t \cos t-2\sin ^2t \cos t=\cos t-4\sin ^2t \cos t=\cos t(1-4\sin ^2t)\]

OpenStudy (anonymous):

thank you!!!

OpenStudy (mertsj):

yw